OpenStudy (anonymous):

What is the derivative of T(p)=87.97+34.96 ln p + 7.91 p^1/2

OpenStudy (anonymous):

I got T'(p)=0 + 34.961 p ln^p-1 + 3.955p^-1/2

OpenStudy (baldymcgee6):

You are going to have to present the original function clearer. Is it T(p)=87.97+34.96*ln(p) + 7.91*p^1/2 ?

OpenStudy (anonymous):

yes

OpenStudy (baldymcgee6):

okay, I am writing up a solution for you :)

OpenStudy (anonymous):

thank you! I think I have it wrong

OpenStudy (baldymcgee6):

http://screencast.com/t/ltqEPHY1G0s

OpenStudy (anonymous):

T'(p)=0+34.96/p+7.91(1.2)p^1.2-1 =34.96/p+9.492p^.2

OpenStudy (baldymcgee6):

^ @gomathi your original function is wrong. Which obviously led you to the wrong derivative

OpenStudy (anonymous):

sorry the last termis P^1/2? then 34.96/p+7.91(1/2)p^1/2-1 34.96/p+3.95p^-1/2 34.96+3.95/p^1/2

OpenStudy (anonymous):

the derivative of any ln is reciprocal of that function

OpenStudy (anonymous):

Thank you! i always forget the rules for ln

OpenStudy (anonymous):

therefore lnp is 1/p

OpenStudy (anonymous):

hope you know constant p0wer of any variable is given by power *p^power minus one so p^1/2 is given by p ^1/2-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hope you follow the rules

OpenStudy (anonymous):

i'll try :)